OpenStudy (anonymous):

How would i find the solution to this equation? log16 32 = x + 2.

OpenStudy (anonymous):

16 is the base

OpenStudy (anonymous):

first we need the number for \[\log_{16}(32)\] i.e. we are trying to solve \[16^y=32\] since \(16=2^4\) and \(32=2^5\) this is the same as solving \[2^{4y}=2^5\] so \[4y=5\] and therefore\[ 9y=\frac{5}{4}\] last job is to solve \[\frac{5}{4}=x+2\]

OpenStudy (anonymous):

typo there, should be \[y=\frac{5}{4}\]